OpenStudy (anonymous):

The area of a triangle is 24 square inches. What is the height of the triangle if the base length is 8 inches? 6 inches 8 inches 12 inches 16 inches

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@aaronq

HanAkoSolo (jamierox4ev3r):

if the area of a triangle is 1/2bh, then we could plug in the values that we already know and solve for height

OpenStudy (anonymous):

1/2 * 8 * 24

HanAkoSolo (jamierox4ev3r):

\(A=\frac{1}{2}bh\) \(24=\frac{1}{2}8 *h\) yep c:

OpenStudy (anonymous):

so its 16

HanAkoSolo (jamierox4ev3r):

i mean...sort of :P you don't multiply all three

HanAkoSolo (jamierox4ev3r):

no you solve this: \(24=\frac{1}{2}8 * h\)

HanAkoSolo (jamierox4ev3r):

its not 16 btw

OpenStudy (anonymous):

half of 8 is 4

OpenStudy (anonymous):

so i multiply 4 by 24

HanAkoSolo (jamierox4ev3r):

right... so you need to divide both sides by 4. \(24=\frac{1}{2}8*h\) 24=4h /4 /4 h=6

HanAkoSolo (jamierox4ev3r):

no you divide 24 by 4. See my process up there? it's like you're solving an equation :)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i need more help on a few mor questions if that is ok (: @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

sure. Just be sure to medal and close this question, since posts get laggy for me after a while :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i tagged you in the next question @Jamierox4ev3r